# From Chaos to Coherence: Effects of High-Order Synaptic Correlations on Neural Dynamics

### v1

@@ Abstract
Real neural networks are known to have high-order cyclic synaptic correlations [11, 16, 24]. We have shown how these connectivity structures influence network dynamics by examining a minimal model of cortical networks designed to capture their essential effects. We show that networks with strong positive correlations have stable or weak chaotic dynamics, which are known to enhance computational performance [47, 48, 49]. Networks with strong negative correlations support robust oscillations, which are found across multiple brain regions (e.g., [50]). Further, the correlation among low-dimensional dynamics, weak chaos, and high-order structures suggests that low-dimensional activity regions in the brain may be highly structured. Furthermore, these findings may also provide insights into the macroscopic statistical properties of other disordered correlated systems, such as spin glasses [51, 52, 53], ecological networks [54, 55], and social networks [56].
Our work establishes a basis for exploring the computational benefits of high-order connectivity structures. Prior studies show that adding low-rank components to random matrices improves performance in discrimination tasks [20] and that specific phase-space patterns relate to learned tasks [57]. Our findings could lead to further improvements and link connectivity to phase-space dynamical motifs. This may illuminate cortical functionality and inspire advances in artificial learning models.
I Supplementary Materials
I.1 Matrix generation
To generate $\alpha$-order cyclic correlations, we start with an $N\times N$ random matrix, where the entries are iid with $\sim\mathcal{N}(0,g^{2}/N)$. Next, we methodically change the sign of the corresponding entries such that the sum of all cycles of length $\alpha$ that end up in that specific entry would be positive with a probability $P$. This probability will determine the strength of the correlations $\rho$, see Fig. S1. We do the same for negative correlations but demand that the sum be negative. The step-by-step algorithm is shown in algorithm 1.
input : $N\times N$ random matrix with $w_{ij}\sim\mathcal{N}(0,g^{2}/N)$
output : $N\times N$ matrix with $\mathrm{Tr}\,W^{\alpha}/N=\rho>0$
1
2Initialize an $N\times N$ random matrix, where the entries are iid with $w_{ij}\sim\mathcal{N}(0,g^{2}/N)$;
3 for $n=\alpha-1:N$ do
4       $\tilde{w}=(w_{n+1,1:n}w_{1:n,1:n}^{\alpha-2})\odot w_{n+1,1:n}$;
5       for $c=1:n$ do
6             if $\tilde{w}_{c,n+1}<0$ then
7                   $w_{c,n+1}\rightarrow-w_{c,n+1}$ by a probability $P$;
8
9             end if
10
11       end for
12
13 end for
Algorithm 1 Generating high-order cyclic correlations
Note that $w_{l,1:k}$ is a $1\times k$ vector, and $\odot$ is the element-wise product.
Figure S1: Numerical value of $\rho$ as a function of the probability $P$, averaged over 30 realizations.
Figure S2: (a) Eigenvalue distribution (black dots) shown for $\rho\approx 0.38$ with $N=6400$. The blue curve depicts the equation of the support $z(\phi)=e^{i\phi}+\rho e^{-2i\phi}$. (b) The same as (a) but with $\rho\approx 0.76$. (c) The observed value of the real part of the rightmost eigenvalue, $\lambda_{1}^{\text{obs}}$ (black) and $g_{\mathrm{eff}}$ (blue) as a function of $N$, with $g=1$ and $\rho=0.76$, averaged over 30 different realizations.
I.2 Numerical details of phase diagram
The probability of observing chaotic, oscillatory, or fixed point behavior was computed by averaging over $300$ different network realizations with $N=1600$. The condition for a fixed-point solution is that the derivative of the trajectories is less than $10^{-4}$ in the final $20\%$ of the total runtime of the dynamics. Otherwise, we estimate the Lyapunov exponents with a cutoff of $0.01$ to separate chaotic from oscillatory solutions. Colors were shaded based on their corresponding probabilities with red, green, and blue, for chaos, oscillations, and fixed points respectively.
Figure S3: Probability of observing chaos for different values of $N$ with (a) $\rho\approx 0$, (b) $\rho\approx 0.23$, and (c) $\rho\approx 0.76$. Data was averaged over $300$ different realizations of the connectivity matrix.
Figure S4: Eigenvalue distributions for $\alpha=3$ (top left), $\alpha=4$ (top right), $\alpha=5$ (bottom left), and $\alpha=6$ (bottom right) with (a) $\rho\approx 0.18$, and (b) $\rho\approx-0.18$. Network size is $N=1600$ and $g_{\mathrm{eff}}\approx 1.1$.
@@ ACKNOWLEDGMENTS
This research was supported in part by grants from NSF NeuroNex (DBI-1707400), NSF (DMS-2235451) and Simons Foundation (MPS-NITMB-00005320) to the NSF-Simons National Institute for Theory and Mathematics in Biology.

### v2

@@ Abstract
Biological neural networks are known to exhibit high-order cyclic synaptic correlations [11, 16, 25]. We have shown how such connectivity structures influence network dynamics by examining a minimal model of cortical networks designed to capture the impact of cyclic correlations. Networks with strong positive correlations support stable or weakly chaotic dynamics, which are known to enhance computational performance [50, 51, 52]. Networks with strong negative correlations support robust oscillations, which are a hallmark of neural dynamics. The strong relation between low-dimensional chaotic dynamics and high-order structures, suggests that low-dimensional neural activity may a consequence of local structure in synaptic connectivity. These findings may also provide insights into the macroscopic statistical properties of other disordered correlated systems, such as spin glasses [53, 54, 55], ecological networks [56, 57], and social networks [58].
Mounting evidence shows that network connectivity patterns shape both learning and computation [59, 44, 21, 60, 61, 62]. Yet, the interplay between the full spectrum of high-order correlations and learning remains largely unexplored. By outlining how higher-order correlations reshape network dynamics, our work provides a concrete framework for testing their computational impact. Pursuing this line of inquiry promises not only to clarify the circuit principles that govern cortical function, but also to inspire new and novel artificial learning models. In addition, embedding these correlations in network models yields biologically plausible architectures that generate specific, testable predictions for future experiments.
@@ ACKNOWLEDGMENTS
This research was supported in part by grants from NSF NeuroNex (DBI-1707400), NIH (R01 MH130416) to NS, XP and KJ, NSF (DMS-2207647) and NSF (DMS-2235451) to KJ, and Simons Foundation (MPS-NITMB-00005320) through the NSF-Simons National Institute for Theory and Mathematics in Biology to NS, XP, KJ, and KEB.